OpenStudy (anonymous):

critical points of f(x)=x-2sqrt(x)

OpenStudy (anonymous):

step1 : find the derivative

OpenStudy (anonymous):

its confusing

OpenStudy (anonymous):

guide me?

OpenStudy (anonymous):

which part ? do it step by step... sure :)

OpenStudy (anonymous):

which part ? do it step by step... sure :)

OpenStudy (anonymous):

which part ? do it step by step... sure :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\(\large f(x) = x - 2 \sqrt{x}\) differentiate both sides, and tell me wat u got

OpenStudy (anonymous):

i dont undesrtand differentiations much

OpenStudy (anonymous):

its easy, just remember below formula :- derivative of \(\large x^n\) is \(\large nx^{n-1}\)

OpenStudy (anonymous):

oh... i am just totally lost. I hope you can still help

OpenStudy (anonymous):

\(\large f(x) = x - 2 \sqrt{x}\) is same as \(\large f(x) = x^1 - 2 x^{\frac{1}{2}}\) differentiate both sides \(\large f'(x) = 1x^{1-1} - 2 \frac{1}{2}x^{\frac{1}{2}-1}\) \(\large f'(x) = 1x^{0} - x^{-\frac{1}{2}}\) \(\large f'(x) = 1 - \frac{1}{\sqrt{x}}\)

OpenStudy (anonymous):

see if you're okay so far btw this question is from 'calculus' right ?

OpenStudy (anonymous):

ya 12th grade

OpenStudy (anonymous):

you got the 1st derivative im not sureabout whats next

OpenStudy (anonymous):

okay good :) next set \(f'(x)\) to 0 and solve x. that gives the cirtical points

OpenStudy (anonymous):

Im confused w/ the fractions :'(

OpenStudy (anonymous):

\(\large 1 - \frac{1}{\sqrt{x}} = 0\) solve \(x\)

OpenStudy (anonymous):

(-1, 0)?

OpenStudy (anonymous):

\(\large 1 - \frac{1}{\sqrt{x}} = 0\) \(\large 1 = \frac{1}{\sqrt{x}} \) \(\large 1 = \sqrt{x} \) \(\large 1 = x \)

OpenStudy (anonymous):

pluggin \(x = 1\) in the given equation :- y = 1 - 2sqrt(1) = 1-2 = -1 thus, the oly critical point for given equation is (1, -1)

OpenStudy (anonymous):

Thanks, Rebecca! :)

OpenStudy (anonymous):

u wlc (: